OpenStudy (amurciano2):

How do I integrate 5^x/ln5

OpenStudy (misty1212):

\[\int b^xdx=\frac{b^x}{ln(b)}\]

OpenStudy (amurciano2):

That helps me with the top part what do I do with the ln5 in the denominator? Do I have to do u-sub?

OpenStudy (amurciano2):

Never mind I get it you place ln5 outside

OpenStudy (empty):

Also it might help to look at this like this: \[5 = e^{\ln 5}\] \[5^x = e^{x\ln 5}\] So really when you take the derivative, that's why you get that extra \(\ln 5\) term.